Consider a hydrogen atom. What wavelength of light (nm) will excite an electron from the n=4 to n=6 principal quantum level


1
Expert's answer
2020-09-21T06:50:49-0400

Rydberg formula for hydrogen atom:

"1/\lambda = R_H(1/n_1^2-1/n_2^2)",

where "\lambda" is the wavelength needed, RH is the Rydberg constant for hydrogen (1.09678*107 m-1), n1 and n2 are principal quantum numbers of the lower and higher energy levels respectively.

"1/\lambda = 1.09678*10^7(1/4^2-1/6^2)"

"\lambda" = 2.626*10-6 m = 2626 nm

Answer: 2626 nm